OpenStudy (anonymous):

what is the y intercept for (-10,-4) and (-1,2)

OpenStudy (solomonzelman):

Want to actually learn ?

OpenStudy (anonymous):

x=5/3, y=0

OpenStudy (shamim):

can u tell me the equation of line which go through (-10,-4) nd (-1,2)

OpenStudy (anonymous):

y1-y2=x1-x2 y=mx+b

OpenStudy (shamim):

no

OpenStudy (anonymous):

y=x(-3/2)+1/2

OpenStudy (anonymous):

wait a while sorry i got it wrong

OpenStudy (anonymous):

y=x(2/3)+8/3

OpenStudy (shamim):

y-y1=((y1-y2)/(x1-x2))(x-x1)